Hello, I am in Physics 102 in college and am currently in a unit on circuits....

Hello,

I am in Physics 102 in college and am currently in a unit on circuits. We are studying voltage, current, resistance, capacitance, etc. The majority of the problems we work on are out of the "Fundamentals of Physics, 9th edition" textbook. I have been using your answers to guide me in my homework, but I would eventually like to be able to have a strategy to conquer these problems on my own. Most of the time I cant figure out the right equations to choose or how to conceptualize these problems. Do you know of any strategies that would allow me to choose the right equations and get my head wrapped around this area?

Thanks!
Answer #1

voltage, current, resistance, capacitance

Okay, These topics are pretty common in Physics.

For you to solve a problem conceptually, make sure to read the given question two or three times to interpret what the question is asking.

There are certain standardized formulas relating to these topics.

V - voltage

I - Current

R - Resistance

C - capacitance

V = IR (Ohm's law)

you can modify this formula to find any unknown provided other two knowns are given.

Then, we have

C = \epsilon A / d

where \epsilon is permittivity and is equal to 8.85e-12 C/m

A is area of a plate of parallel plate capacitor

d is the distance between plates.

Here, you can find value of capacitance.

Sometimes, you may get questions are capacitors in series or parallel and resistors in series or parallel.

For capacitors in series,

1/C = 1/C1 + 1/C2 ----- so on

For capacitors in parallel,

C = C1 + C2 --------------- so on

For resistors in series,

R = R1 + R2 ------------- so on

For resistors in parallel,

1/R = 1/R1 + 1/R2------------ so on

One thing which I would like to add here is practice as much as you can. In this way, you will learn how to approach a problem.

Let me know if you need more information.